Solve the initial value problem

y′′+4y′+4y=0, y(−1)=4, y′(−1)=4
1
Expert's answer
2020-07-16T19:36:04-0400

"y''+4y'+4y=0"

initial values:

"y(-1)=4"

"y'(-1)=4"


Step by step solution: "\tfrac{d^2y(x)}{dx^2} + 4 \tfrac{dy(x)}{dx} + 4y(x) = 0"

1-step: Assume a solution will be proportioanl to "e^{\lambda x}" for some constant "\lambda" . Substitute "y(x) = e^{\lambda x}" into the differential equation:


"\tfrac{d^2 (e^{\lambda x})}{dx^2} + 4 \tfrac{d (e^{\lambda x})}{dx} + 4e^{\lambda x} = 0"


2-step: Substiute "\tfrac{d^2 (e^{\lambda x})}{dx^2} = \lambda ^2 e^{\lambda x}" and "\tfrac{d (e^{\lambda x})}{dx} = \lambda e^{\lambda x}"


"\lambda^2 e^{\lambda x} + 4\lambda e^{\lambda x} + 4e^{\lambda x} = 0"


3-step: Factor out "e^{\lambda x}"


"(\lambda^2+4\lambda+4)e^{\lambda x} = 0"


4-step: Since "e^{\\lambda x} {=}\\mathllap{\/\\,}\t0" for any finite "\lambda" , the zeros must come from the polynomial "\lambda^2+4\lambda+4=0"

5-step: Factor


"(\lambda + 2)^2 = 0"


6-step: Solve for "\lambda"


"\lambda = -2" or "\lambda = -2"


7-step: The multiplicity of the root "\lambda = -2" is 2 which gives y1(x) = C1 "e^{-2x}" , y2(x) = C2 "e^{-2x} x" as solutions, where C1 and C2 are arbitrary constants.

The general solution is the sum of the above solutions:


y(x) = y1(x) + y2(x) = C1"e^{-2x}" + C2"e^{-2x}x"


8-step: Solve for the unknown constants using the given initial conditions

Compute "\tfrac{dy(x)}{dx}:"


"\tfrac{dy(x)}{dx} = \tfrac{d }{dx}" (C1"e^{-2x} +" C2"e^{-2x}x" ) = - 2C1"e^{-2x} +" C2"e^{-2x} -" 2C2"e^{-2x} x"



9-step: Substitute y(-1) = 4 into y(x) = C1"e^{-2x} + e^{-2x}x"C2


"e^{2}" C1 - "e^{2}" C2 = 4


10-step: Substitute y'(-1) = 4 into "\tfrac{dy(x)}{dx} = -2e^{-2x}" C1 + "e^{-2x}" C2 - "2e^{-2x}x" C2


-2C1"e^{2} + 3e^{2}" C2 = 4


11-step: Solve the system:


C1 = "\tfrac{16}{e^{2}}"


C2 = "\tfrac{12}{e^{2}}"


Substitute C1 = "\tfrac{16}{e^{2}}" and C2 = "\tfrac{12}{e^{2}}" into y(x) = "e^{-2x}" C1 + "e^{-2x}x" C2


Answer:"y(x) = 4e^{-2(x+1)}(3x+4)"